OpenStudy (anonymous):

What is the solution to the equation 2^(1+2x)= 3 ?

OpenStudy (anonymous):

@phi @experimentX @Luis_Rivera

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

use logarithms

OpenStudy (anonymous):

i don't know how to

OpenStudy (anonymous):

log to base 10 log2^(1+2x)=log3 apply log properties (1+2x)log2=log3 now some algebra 1+2x=log3/log2 2x=(log3/log2)-1 x=((log3/log2)-1)/2,do the calculations for the log on calculator and use the original equation to see if your answer is right

OpenStudy (anonymous):

i no has calculator

OpenStudy (anonymous):

can you do it for me?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

that's not one of the choices

OpenStudy (anonymous):

pliz get a calculator and read something about log thanks

OpenStudy (anonymous):

x ≈ –0.292 x ≈ –0.185 x ≈ 0.185 x ≈ 0.292

OpenStudy (anonymous):

@pasta you did it wrong

OpenStudy (anonymous):

got you!!! so you have a calculator.thats the answer x=0.2924812503606

OpenStudy (anonymous):

oh you ;)